# Formation of GW230529 from Isolated Binary Evolution and Its Electromagnetic Counterparts

### v1

@@ Abstract
In this Letter, we explore the formation of the mass-gap black hole-neutron star (mgBHNS) merger detected in gravitational wave (GW) event, i.e., GW230529, from the isolated binary evolution channel, and study potential signatures of its electromagnetic signals. By adopting the ‘delayed’ supernova prescription and reasonable model realizations, our population synthesis simulation results can simultaneously match the inferred event rate densities of GW230529-like mgBHNS and total BHNS mergers, as well as the population distribution of the BH mass in BHNS mergers reported by the LIGO-Virgo-KAGRA Collaboration. Thus, we conclude that the recently-discovered mgBHNS merger, GW230529, can be explained through the isolated binary evolution channel. Considering the equation of states of AP4 and DD2, the probabilities that GW230529 can make tidal disruption are $12.8\%$ and $63.2\%$, respectively. If GW230529 is a disrupted event, the associated kilonova is predicted to have an apparent magnitude of $\sim 23-24\,{\rm{mag}}$, and hence, can be detected by the present survey projects and LSST. Since GW230529 could be an off-axis event inferred from the GW observation, its associated gamma-ray burst (GRB) might be too dim to be observed by $\gamma$-ray detectors, interpreting the lack of GRB observations. The detection of GW230529 confirms the existence of mgBHNS mergers formed through the isolated binary evolution channel, suggesting that BHNS mergers are still likely to be multimessenger sources that emit GWs, GRBs, and kilonovae. Although mgBHNS mergers account for $\sim 60\%$ cosmological BHNS population, we find that $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers.
@@
software: COMPAS, (Team COMPAS: Riley, J. et al., 2022); Python, https://www.python.org; Matlab, https://www.mathworks.com; OriginPro, https://www.originlab.com/originpro
@@ 1 Introduction
Black hole–neutron star (BHNS) mergers are prime search targets for the ground-based gravitational-wave (GW) detectors, including LIGO (LIGO Scientific Collaboration et al., 2015), Virgo (Acernese et al., 2015), and KAGRA (Aso et al., 2013). Until the end of the GW third observing run (O3), the first two BHNS mergers were identified by the LIGO-Virgo-KAGRA (LVK) Collaboration (Abbott et al., 2021a), with GW200105 being a merger between a $8.9^{+1.2}_{-1.5}M_{\odot}$ BH and a $1.9^{+0.3}_{-0.2}M_{\odot}$ NS, and GW200115 being a merger between a $5.7^{+1.8}_{-2.1}M_{\odot}$ BH and a $1.5^{+0.7}_{-0.3}M_{\odot}$ NS. During O3, the LVK Collaboration reported three additional marginal BHNS candidates (Abbott et al., 2021b; The LIGO Scientific Collaboration et al., 2021; Abbott et al., 2023), with GW190426_152155 involving a $5.7^{+3.9}_{-2.3}M_{\odot}$ BH and a $1.5^{+0.8}_{-0.5}M_{\odot}$ NS, GW190917_114630 involving a $9.7^{+3.4}_{-3.9}M_{\odot}$ BH and a $2.1^{+1.1}_{-0.4}M_{\odot}$ NS, as well as GW191219_163120 involving a $31.1^{+2.2}_{-2.8}M_{\odot}$ BH and a $1.17^{+0.07}_{-0.06}M_{\odot}$ NS. GWTC-3 also includes a puzzling GW event, GW190814, composed of a $22.2-24.3\,M_{\odot}$ primary BH and a $2.50-2.67\,M_{\odot}$ secondary compact object (Abbott et al., 2020). The origin of its secondary object is unclear, which could be either a low-mass BH or a massive NS. The LVK Collaboration also discovered a similar, but marginal GW candidate, namely GW200210_092254 (Abbott et al., 2023), with component masses inferred to be $24.1^{+7.5}_{-4.6}M_{\odot}$ and $2.83^{+0.47}_{-0.42}M_{\odot}$, respectively.
@@
Binary NSs (BNSs) and BHNSs have long been proposed to be progenitors of some fast-evolving EM transients, including gamma-ray bursts (GRBs; Paczynski, 1986, 1991; Eichler et al., 1989; Narayan et al., 1992; Zhang, 2018; Gottlieb et al., 2023) and kilonovae (Li & Paczyński, 1998; Metzger et al., 2010). While BNS mergers were confirmed as the origin of GRBs and kilonovae thanks to the associations between GW170817 (Abbott et al., 2017a), GRB170817A (Abbott et al., 2017b; Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018), and AT2017gfo (e.g., Abbott et al., 2017c; Arcavi et al., 2017; Coulter et al., 2017; Drout et al., 2017; Evans et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Kilpatrick et al., 2017; Pian et al., 2017; Smartt et al., 2017), the joint observations of the O3 BHNS GW candidates and their associated EM counterparts, especially for kilonova emissions, were absent. One possible reason for the lack of kilonova detection following O3 BHNS GW signals could be the challenge of rapidly achieving full distance and volumetric coverage for the probability maps of the LVK Collaboration within the short kilonova duration by current survey projects (e.g., Coughlin et al., 2020; Kasliwal et al., 2020; Gompertz et al., 2020; Anand et al., 2021). Unlike BNS mergers, which typically eject a certain amount of materials to produce EM counterparts, the NS components in some merging BHNS binaries may directly plunge into their BH companions without generating any observable EM signals. NS tidal disruption tends to occur if the BHNS binaries have a low-mass BH with a high orbital aligned spin and a low-mass NS with a stiff EoS (Shibata & Taniguchi, 2011; Foucart, 2012; Kyutoku et al., 2013, 2015; Kawaguchi et al., 2016; Foucart et al., 2018; Zhu et al., 2020, 2021b; Hayashi et al., 2021; Sarin et al., 2022). More specifically, since primary BHs (produced from initially more massive star) formed through the classical common envelope scenario typically possess near-zero aligned spins (Qin et al., 2018; Fuller et al., 2019; Belczynski et al., 2020), tidal disruptions of most cosmological BHNS mergers are expected to occur if the BHs have a mass of $\lesssim 6-7\,M_{\odot}$ and the NSs have a mass of $\lesssim 1.5\,M_{\odot}$ (e.g., Zhu et al., 2021a, 2022b). If the mass gap does exist, the mass space that allows NS tidal disruption and produces bright EM signals could be limited. Because O3 BHNS candidates have BH masses of $\gtrsim 5\,M_{\odot}$ inferred from GW observations, their posterior mass distributions mostly lie outside the tidal disruption mass range (Abbott et al., 2021a; Zhu et al., 2021a, 2022b; D’Orazio et al., 2022). Thus, it is likely that the EM counterparts associated with these BHNS candidates were intrinsically missing. Additionally, although a few O3 BHNS candidates still have low probabilities of undergoing tidal disruption and producing kilonova emissions, the brightness of these kilonovae might be too dim to be detected by current survey telescopes due to their remote distance from us (Zhu et al., 2021a).
@@
The discovery of mgBHNS mergers through GW observations can provide an opportunity to constrain SN mechanisms and isolated formation channel of compact binaries. Furthermore, mgBHNS mergers are more likely to have tidal disruption and produce observable EM counterparts. Multimessenger observations of BHNS mergers between GWs and EMs can be more easily achieved if mgBHNS mergers exist. Most recently, the LVK Collaboration reported the first mgBHNS merger, GW230529, detected in the first part of the fourth observing run (The LIGO Scientific Collaboration et al., 2024). By adopting the combined posterior results inferred by the low-spin prior of the secondary component (The LIGO Scientific Collaboration et al., 2024), the primary BH mass of GW230529 is $M_{\rm BH}=3.6^{+0.7}_{-1.2}M_{\odot}$ nearly filling in the mass gap, while its secondary NS mass is $M_{\rm NS}=1.43^{+0.59}_{-0.19}M_{\odot}$. The effective inspiral-spin and effective precessing-spin of GW230529 are $\chi_{\rm eff}=-0.10^{+0.10}_{-0.17}$ and $\chi_{\rm p}=0.40^{+0.37}_{-0.34}$, respectively. The source has a redshift of $z=0.043^{+0.023}_{-0.021}$, corresponding to a luminosity distance of $D_{\rm L}=197^{+107}_{-96}{\rm Mpc}$. The LIGO Scientific Collaboration et al. (2024) gave the rate density of $\mathcal{R}_{230529}=55^{+127}_{-47}{\rm Gpc}^{-3}{\rm yr}^{-1}$ for compact object coalescences with properties similar to GW230529. Similar to the estimation of $\mathcal{R}_{230529}$, they also updated a total BHNS rate density to $\mathcal{R}_{\rm BHNS}=85^{+116}_{-57}{\rm Gpc}^{-3}{\rm yr}^{-1}$ by including the other two O3 BHNS mergers (GW200105 and GW200115) [^1] By considering that GW200105, GW200115, and GW230529 belong to a single population class, as well as taking into account some marginal events detected before GW230529, The LIGO Scientific Collaboration et al. (2024) also gave another BHNS rate density of $\mathcal{R}_{\rm BHNS}=94^{+109}_{-64}{\rm Gpc}^{-3}{\rm yr}^{-1}$.. In this Letter, we explore the formation of GW230529 through the isolated binary evolution channel and study the properties of its associated EM counterparts.
@@ 2.1 Method
In this section, we employ the rapid binary population synthesis code COMPAS (Stevenson et al., 2017; Vigna-Gómez et al., 2018; Neijssel et al., 2019; Team COMPAS: Riley, J. et al., 2022) to explore the formation of GW230529 through the isolated binary evolution channel. To characterize star formation environments, we adopt a metallicity-specific star formation rate density model suggested in Broekgaarden & Berger (2021) which combines the Madau & Dickinson (2014) star formation density with the Panter et al. (2004) mass-metallicity relation. To determine compact object masses during core-collapse SNe of stars, the ‘delayed’ SN prescription (Fryer et al., 2012) is adopted, which allows for the production of mass-gap BHs (see also Mandel & Müller, 2020). Based on the metallicity-specific star formation rate density prescription, we generate and evolve binary systems according to the fiducial population synthesis model described in Table 1 of Broekgaarden et al. (2021), with the maximum NS mass changed to $M_{\rm NS,max}=2.2\,M_{\odot}$, and the velocity dispersion of BH natal kick, modeled with a Maxwellian distribution, to $\sigma_{\rm BH}=100\,{\rm km}\,{\rm s}^{-1}$ following with the measurements on the proper motions of Galactic BH X-ray binaries (Atri et al., 2019).
@@ 2.2 Results
Model $\mathcal{R}_{230529}$ $\mathcal{R}_{\rm BHNS}$ CI
@@
$\alpha 1\sigma 100$ $65$ $180$ 0.64
$\alpha 2\sigma 100$ $58$ $146$ 0.38
$\alpha 5\sigma 100$ $47$ $74$ 0.09
$\alpha 10\sigma 100$ $54$ $71$ 0.13
$\alpha 1\sigma 265$ $41$ $134$ 0.29
$\alpha 2\sigma 265$ $49$ $115$ 0.09
$\alpha 5\sigma 265$ $37$ $60$ 0.36
$\alpha 10\sigma 265$ $34$ $52$ 0.50
Note. — The columns are (1) the population synthesis model; (2) the simulated local merger rate density of GW230529-like events in ${\rm Gpc}^{-3}\,{\rm yr}^{-1}$; (3) the simulated local merger rate density of total BHNS mergers in ${\rm Gpc}^{-3}\,{\rm yr}^{-1}$; (4) the credible interval.
@@
The simulated local event rate densities of GW230529-like events and BHNS mergers are listed in Table 1 and are displayed in Figure 1. Based on the probability density functions of $\mathcal{R}_{230529}$ and $\mathcal{R}_{\rm BHNS}$, assuming they are independent, we calculate the credible interval of our population synthesis simulations for the LVK’s observations. The event rate densities of our population synthesis simulations are all within the $60\%$ credible intervals of observed $\mathcal{R}_{230529}$ and $\mathcal{R}_{\rm BHNS}$. We find that the models with a lower common-envelope efficiency ($\alpha_{\rm CE}\lesssim 2$) indicates a better agreement with the LVK’s observations when the natal kick is strong ($\sigma_{\rm NS}=265\,{\rm km}\,{\rm s}^{-1}$), whereas those with higher common-envelope efficiency ($\alpha_{\rm CE}\gtrsim 5$) suggests better consistency with the observations when the kick velocity is relatively low ($\sigma_{\rm NS}=100\,{\rm km}\,{\rm s}^{-1}$). Among these population synthesis models, the models of $\alpha 5\sigma 100$ and $\alpha 2\sigma 265$, exhibit credible intervals lower than 10% within the GW observations.
@@
Figure 2: NS and BH masses of our population synthesis simulations. The grey dotted lines represent the credible interval regions of the simulation results, spanning from $10\%$ to $90\%$, estimated using kernel density estimation. The medians (star points) along with their corresponding $90\%$ credible intervals (solid lines) are illustrated for the posterior distributions of GW200105 (green), GW200115 (blue), and GW230529 (orange). The right panel shows the population distribution of the BH mass inferred by the GW observations (orange histogram; The LIGO Scientific Collaboration et al., 2024), contrasted with that derived from our population synthesis simulations (grey histogram). For an EoS of AP4 (DD2), non-spinning BHNS mergers with component masses located at the bottom left parameter space of the pink (light pink) dashed line can allow tidal disruption.
@@
By combining the population synthesis simulation results of $\alpha 5\sigma 100$ and $\alpha 2\sigma 265$, we show the distribution of the BH mass and NS mass for simulated BHNS population in Figure 2. The medians with the $90\%$ credible intervals of the inferred posterior samples for GW200115, GW200105, and GW230529 are also displayed. One can find that the combined model predicts that the majority of BHNS mergers can have a BH mass of $2.2-11\,M_{\odot}$, and a NS mass of $1.25-2.2\,M_{\odot}$, with a concentration of $\sim 1.3\,M_{\odot}$. The inferred properties of GW200105, GW200115, and GW230529 lie within the $90\%$ credible interval of our simulated BHNS population. In particular, the medians with a large fraction of the posterior for the mgBHNS event, i.e., GW230529, recently reported by the LVK Collaboration significantly overlap with the highest probability region of the simulated BHNS population. We find that the predicted BH mass distribution is basically consistent with that of GW BHNS population inferred by the presently detected three high-confidence BHNS mergers, particularly for the mass space of $\gtrsim 5\,M_{\odot}$. Compared with the present GW observation, where the BH mass peaks at $\sim 4\,M_{\odot}$, our simulated BH population has a lower mass peak of $3.2\,M_{\odot}$.
@@
Overall, our population synthesis simulation results can simultaneously match the inferred evet rate densities of GW230529-like mgBHNS and total BHNS mergers, along with the GW population distribution of the BH mass in BHNS mergers. The posterior masses of GW230529 given by the LVK Collaboration are close to the highest probability region of the simulated BHNS population. Therefore, one can expect that GW230529 can likely originate from the isolated binary evolution channel.

### v2

@@ Abstract
In this Letter, we explore the formation of the mass-gap black hole-neutron star (mgBHNS) merger detected in gravitational wave (GW) event, i.e., GW230529, from the isolated binary evolution channel, and study potential signatures of its electromagnetic counterparts. By adopting the ‘delayed’ supernova prescription and reasonable model realizations, our population synthesis simulation results can simultaneously match the rate densities of mgBHNS and total BHNS mergers inferred from the population analyses, along with the population distribution of the BH mass in BHNS mergers reported by the LIGO-Virgo-KAGRA Collaboration. Because GW230529 contributes significantly to the inferred mgBHNS rate densities, we suggest that GW230529 can be explained through the isolated binary evolution channel. Considering the AP4 (DD2) equation of state, the probability that GW230529 can make tidal disruption is $12.8\%$ ($63.2\%$). If GW230529 is a disrupted event, its kilonova peak apparent magnitude is predicted $\sim 23-24\,{\rm{mag}}$, and hence, can be detected by the present survey projects and LSST. Since GW230529 could be an off-axis event inferred from the GW observation, its associated gamma-ray burst (GRB) might be too dim to be observed by $\gamma$-ray detectors, interpreting the lack of GRB observations. Our study suggests the existence of mgBHNS mergers formed through the isolated binary evolution channel due to the discovery of GW230529, indicating that BHNS mergers are still likely to be multimessenger sources that emit GWs, GRBs, and kilonovae. Although mgBHNS mergers account for $\sim 50\%$ cosmological BHNS population, we find that $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers.
@@
software: COMPAS (version 02.39.00; Team COMPAS: Riley, J. et al., 2022); GWOSC, https://gwosc.org; Python, https://www.python.org; Matlab, https://www.mathworks.com; OriginPro, https://www.originlab.com/originpro
@@ 1 Introduction
Black hole–neutron star (BHNS) mergers are prime search targets for the ground-based gravitational-wave (GW) detectors, including LIGO (LIGO Scientific Collaboration et al., 2015), Virgo (Acernese et al., 2015), and KAGRA (Aso et al., 2013). Until the end of the GW third observing run (O3), the first two BHNS mergers were identified by the LIGO-Virgo-KAGRA (LVK) Collaboration (Abbott et al., 2021a), with GW200105 being a merger between a $8.9^{+1.2}_{-1.5}M_{\odot}$ BH and a $1.9^{+0.3}_{-0.2}M_{\odot}$ NS, and GW200115 being a merger between a $5.7^{+1.8}_{-2.1}M_{\odot}$ BH and a $1.5^{+0.7}_{-0.3}M_{\odot}$ NS (all measurements quoted at the 90% credible level). During O3, the LVK Collaboration reported three additional marginal BHNS candidates (Abbott et al., 2021b; The LIGO Scientific Collaboration et al., 2021; Abbott et al., 2023), with GW190426_152155 involving a $5.7^{+3.9}_{-2.3}M_{\odot}$ BH and a $1.5^{+0.8}_{-0.5}M_{\odot}$ NS, GW190917_114630 involving a $9.7^{+3.4}_{-3.9}M_{\odot}$ BH and a $2.1^{+1.1}_{-0.4}M_{\odot}$ NS, as well as GW191219_163120 involving a $31.1^{+2.2}_{-2.8}M_{\odot}$ BH and a $1.17^{+0.07}_{-0.06}M_{\odot}$ NS. GWTC-3 also includes a puzzling GW event, GW190814, composed of a $22.2-24.3\,M_{\odot}$ primary BH and a $2.50-2.67\,M_{\odot}$ secondary compact object with unclear origin (Abbott et al., 2020). The LVK Collaboration also discovered a similar, but marginal GW candidate, namely GW200210_092254 (Abbott et al., 2023), with component masses inferred to be $24.1^{+7.5}_{-4.6}M_{\odot}$ and $2.83^{+0.47}_{-0.42}M_{\odot}$, respectively.
@@
Binary NSs (BNSs) and BHNSs have long been proposed to be progenitors of some fast-evolving EM transients, including gamma-ray bursts (GRBs; Paczynski, 1986, 1991; Eichler et al., 1989; Narayan et al., 1992; Zhang, 2018; Gottlieb et al., 2023) and kilonovae (Li & Paczyński, 1998; Metzger et al., 2010). While BNS mergers were confirmed as the origin of GRBs and kilonovae thanks to the associations between GW170817 (Abbott et al., 2017a), GRB170817A (Abbott et al., 2017b; Goldstein et al., 2017; Savchenko et al., 2017; Zhang et al., 2018), and AT2017gfo (e.g., Abbott et al., 2017c; Arcavi et al., 2017; Coulter et al., 2017; Drout et al., 2017; Evans et al., 2017; Kasen et al., 2017; Kasliwal et al., 2017; Kilpatrick et al., 2017; Pian et al., 2017; Smartt et al., 2017), the joint observations of the O3 BHNS GW candidates and their associated EM counterparts, especially for kilonova emissions, were absent. One possible reason for the lack of kilonova detection following O3 BHNS GW signals could be the challenge of rapidly achieving full distance and volumetric coverage for the probability maps of the LVK Collaboration within the short kilonova duration by current survey projects (e.g., Coughlin et al., 2020; Kasliwal et al., 2020; Gompertz et al., 2020; Anand et al., 2021). Unlike BNS mergers, which typically eject a certain amount of materials to produce EM counterparts, the NS components in some merging BHNS binaries may directly plunge into their BH companions without generating any observable EM signals. NS tidal disruption tends to occur if the BHNS binaries have a low-mass BH with a high orbital aligned spin and a low-mass NS with a stiff EoS (Shibata & Taniguchi, 2011; Foucart, 2012; Kyutoku et al., 2013, 2015; Kawaguchi et al., 2016; Foucart et al., 2018; Zhu et al., 2020, 2021b; Hayashi et al., 2021; Sarin et al., 2022). More specifically, since primary BHs (produced from initially more massive star) formed through the classical CE scenario typically possess near-zero aligned spins (Qin et al., 2018; Fuller et al., 2019; Belczynski et al., 2020), tidal disruptions of most cosmological BHNS mergers are expected to occur if the BHs have a mass of $\lesssim 6-7\,M_{\odot}$ and the NSs have a mass of $\lesssim 1.5\,M_{\odot}$ (e.g., Zhu et al., 2021a, 2022b). If the mass gap does exist, the mass space that allows NS tidal disruption and produces bright EM signals could be limited. Because O3 BHNS candidates have BH masses of $\gtrsim 5\,M_{\odot}$ inferred from GW observations, their posterior mass distributions mostly lie outside the tidal disruption mass range (Abbott et al., 2021a; Zhu et al., 2021a, 2022b; D’Orazio et al., 2022). Thus, it is likely that the EM counterparts associated with these BHNS candidates were intrinsically missing. Additionally, although a few O3 BHNS candidates still have low probabilities of undergoing tidal disruption and producing kilonova emissions, the brightness of these kilonovae might be too dim to be detected by current survey telescopes due to their remote distance from us (Zhu et al., 2021a).
@@
The discovery of mgBHNS mergers through GW observations can provide an opportunity to constrain SN mechanisms and isolated formation channel of compact binaries. Furthermore, mgBHNS mergers are more likely to have tidal disruption and produce observable EM counterparts. Multimessenger observations of BHNS mergers between GWs and EMs can be more easily achieved if mgBHNS mergers exist. Most recently, the LVK Collaboration reported the first mgBHNS merger, GW230529, detected in the first part of the fourth observing run (The LIGO Scientific Collaboration et al., 2024). By adopting the combined posterior results inferred by the low-spin prior of the secondary component (i.e., the posterior sample of Combined_PHM_lowSecondarySpin; The LIGO Scientific Collaboration et al., 2024), , the primary BH mass of GW230529 is $M_{\rm BH}=3.6^{+0.7}_{-1.2}M_{\odot}$ nearly filling in the mass gap, while its secondary NS mass is $M_{\rm NS}=1.43^{+0.59}_{-0.19}M_{\odot}$. The effective inspiral-spin and effective precessing-spin of GW230529 are $\chi_{\rm eff}=-0.10^{+0.10}_{-0.17}$ and $\chi_{\rm p}=0.40^{+0.37}_{-0.34}$, respectively. The source has a redshift of $z=0.043^{+0.023}_{-0.021}$, corresponding to a luminosity distance of $D_{\rm L}=197^{+107}_{-96}{\rm Mpc}$. Based on the population properties of currently detected BHNS mergers using the NSBH-pop model reported by The LIGO Scientific Collaboration et al. (2024), we find the rate densities of mgBHNS mergers with a primary BH mass $\lesssim 5\,M_{\odot}$ and total BHNS mergers to be $R_{\rm mgBHNS}=18^{+58}_{-16}\,{\rm Gpc}^{-3}{\rm yr}^{-1}$ and $R_{\rm BHNS}=40^{+77}_{-29}\,{\rm Gpc}^{-3}{\rm yr}^{-1}$, respectively. Without the inclusion of GW230529, these rate densities change to be $R_{\rm mgBHNS}=0^{+10}_{-0}\,{\rm Gpc}^{-3}{\rm yr}^{-1}$ and $R_{\rm BHNS}=18^{+41}_{-15}\,{\rm Gpc}^{-3}{\rm yr}^{-1}$. Thus, GW230529 significantly increases the inferred rate of mgBHNS mergers.
In this Letter, we explore the formation of GW230529 through the isolated binary evolution channel and study the properties of its associated EM counterparts. Hereafter, the combined posterior sample inferred by the default low-spin prior of the secondary component released by the LVK Collaboration is employed for our studies on individual GW sources, including GW200105, GW200115, and GW230529 (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024).
@@ 2.1 Method
In this section, we employ the rapid binary population synthesis code COMPAS (version 02.39.00; Stevenson et al., 2017; Vigna-Gómez et al., 2018; Neijssel et al., 2019; Team COMPAS: Riley, J. et al., 2022) to explore the formation of GW230529 through the isolated binary evolution channel. To determine compact object masses during core-collapse SNe (CCSNe) of stars, the ‘delayed’ SN prescription (Fryer et al., 2012) is adopted, which allows for the production of mass-gap BHs (see also the stochastic recipe developed by Mandel & Müller, 2020). We generate and evolve binary systems according to the fiducial population synthesis model described in Table 1 of Broekgaarden et al. (2021), with the maximum NS mass changed to $M_{\rm NS,max}=2.2\,M_{\odot}$. We refer readers to Table 4 in Appendix for further details on other initial conditions, parameter settings, and simulation settings we used.
@@ 2.2 Results
Model $\mathcal{R}_{\rm mgBHNS}$ $\mathcal{R}_{\rm BHNS}$ CI
@@
$\alpha 1\sigma 100$ $57$ $149$ 0.89
$\alpha 2\sigma 100$ $93$ $154$ 0.84
$\alpha 5\sigma 100$ $74$ $86$ 0.42
$\alpha 10\sigma 100$ $58$ $67$ 0.47
$\alpha 1\sigma 265$ $9$ $39$ 0.62
$\alpha 2\sigma 265$ $17$ $37$ 0.16
$\alpha 5\sigma 265$ $22$ $28$ 0.41
$\alpha 10\sigma 265$ $16$ $25$ 0.30
Note. — The columns are (1) the population synthesis model; (2) the simulated local merger rate density of mgBHNS events in ${\rm Gpc}^{-3}\,{\rm yr}^{-1}$; (3) the simulated local merger rate density of total BHNS mergers in ${\rm Gpc}^{-3}\,{\rm yr}^{-1}$; (4) the credible interval.
@@
The simulated local event rate densities of mgBHNS and BHNS mergers are listed in Table 1 and are displayed in Figure 1. In Table 1, we also calculate the confidence intervals for our simulated population synthesis rate densities to fall within the posterior distribution of $\mathcal{R}_{\rm mgBHNS}$ and $\mathcal{R}_{\rm BHNS}$ obtained from the population analyses reported by the LVK Collaboration (The LIGO Scientific Collaboration et al., 2024). The event rate densities of our population synthesis simulations are all within the $90\%$ credible intervals of observed $\mathcal{R}_{\rm mgBHNS}$ and $\mathcal{R}_{\rm BHNS}$. We find that the models with a higher SN kick indicate a better agreement with the LVK’s observations. Among these population synthesis models, the model of $\alpha 2\sigma 265$ exhibits the best consistency with the GW observations.
@@
Figure 2: NS and BH masses of our population synthesis simulations. The grey dotted lines represent the credible interval regions of the simulation results, spanning from $10\%$ to $90\%$, estimated using kernel density estimation. The source-frame medians (star points) along with their corresponding $90\%$ credible intervals (solid lines) are illustrated for the posterior distributions of GW200105 (green), GW200115 (blue), and GW230529 (orange). The right panel shows the population distribution of the BH mass inferred by the GW observations (orange histogram; The LIGO Scientific Collaboration et al., 2024), contrasted with that derived from our population synthesis simulations (grey histogram). For an EoS of AP4 (DD2), non-spinning BHNS mergers with component masses located at the bottom left parameter space of the pink (light pink) dashed line can allow tidal disruption.
@@
We show the distribution of the BH mass and NS mass for the $\alpha 2\sigma 265$ population model in Figure 2. The source-frame medians with the $90\%$ credible intervals of the inferred posterior samples for GW200105, GW200115, and GW230529 are also displayed. One can find that the combined model predicts that the majority of BHNS mergers can have a BH mass of $2.2-13.5\,M_{\odot}$, and a NS mass of $1.25-2.2\,M_{\odot}$, with a concentration of $\sim 1.3\,M_{\odot}$. The inferred properties of GW200105, GW200115, and GW230529 lie within the $70\%$ credible interval of our simulated BHNS population. In particular, the medians with a large fraction of the posterior for the mgBHNS event, i.e., GW230529, recently reported by the LVK Collaboration significantly overlap with the highest probability region of the simulated BHNS population. We find that the predicted BH mass distribution is basically consistent with that of GW BHNS population inferred by the presently detected three high-confidence BHNS mergers. Compared with the current GW observations, where the BH mass peaks at $\sim 4\,M_{\odot}$, our simulated BH population has a lower mass peak of $3.4\,M_{\odot}$, which is approximately at the median of the posterior BH mass distribution of GW230529.
@@
Overall, the posterior masses of GW230529 given by the LVK Collaboration are close to the highest probability region of the simulated BHNS population. Our population synthesis simulation results can simultaneously match the inferred event rate densities of mgBHNS and total BHNS mergers, along with the GW population distribution of the BH mass in BHNS mergers. Because GW230529 contributes significantly to the inferred mgBHNS rate densities, one can expect that GW230529 can likely originate from the isolated binary evolution channel.